FIND THE BALANCE FOR A FOR P DOLLARS INVSTED AT RATE r for t years, compounded quarterly
p=$1000
r=6% compounded continuously

Answer by stanbon(75887)

FIND THE BALANCE FOR A FOR P DOLLARS INVSTED AT RATE r for t years, compounded quarterly
p=$1000
r=6% compounded continuously
----------------------
A(t) = P(1+(r/n))^(nt)
A(t) = 1000(1+(0.6/4))^(4t)
A(t) = 1000(1.15)^(4t)
==============================
Cheers,
Stan H.